# Challenges and recommendations for Electronic Health Records data extraction and preparation for dynamic prediction modelling in hospitalized patients - a practical guide

### v1

@@ 5 Discussion and Conclusion
Adhering to the “garbage in, garbage out” principle, prediction models rely heavily on the quality and relevance of the input data to generate meaningful and reliable predictions. Data extraction and preparation for predictive modelling using EHR data are resource-intensive processes, with time and cost varying depending on the maturity of the extraction framework, the prediction task and the team’s experience in working with EHR data. Data quality often varies significantly across EHR systems and extraction processes, as noted by Weiskopf et al. (Weiskopf and Weng, 2013), and so does data preparation too. Issues such as data gaps, temporal leaks, incorrect linking, inconsistencies in clinical concept and terminology mapping can affect the quality of extracted datasets and compromise the model’s performance and applicability. To address such challenges, we proposed a list of practical recommendations informed by our experiences with EHR data and insights from published studies. We organized the challenges and recommendations into: cohort definition, outcome definition, feature engineering and data cleaning; the first three categories can also be consulted when planning a project. A clear definition of the prediction task or research question, of the intended use and the intended users of a prediction model are a first critical step for defining the outcome, the cohort and the features of interest (Arbet et al., 2021; Honeyford et al., 2022). It is not uncommon to deem the EHR data inadequate for the prediction task, before proceeding with the model building phase (Thuraisingam et al., 2021).
@@
For cohort definition, we recommend extracting a broader patient context beyond the immediate focus, assessing the completeness of data used for inclusion/exclusion criteria and its availability at prediction time, and careful definition of episodes of interest for prediction. Outcomes can be derived in different ways, all with advantages and shortcomings. We generally advise against the use of ICD codes, unless carefully assessed as appropriate for the prediction task. Good understanding of hospital’s processes, thorough verification of outcomes derived in code, manual inspection of labels and agreement between data sources can also prevent incorrect outcome definition. Mapping terminology to coding systems (e.g., ICD, CPT, LOINC, SNOMED-CT) can facilitate feature engineering. However, not all items in the EHR system are aligned with standardized terminologies (e.g., LOINC codes for lab results are often not used). Implementing terminology mapping during the data extraction phase (e.g., through OMOP CDM) can significantly reduce the effort required during data preparation. High-quality and completeness of data extraction documentation and good understanding of the underlying clinical concepts and healthcare processes are essential to support feature engineering. For feature engineering in prediction settings, the timestamp when a clinical item is available in the system is of most interest, as this is the time at which predictor values become available for prediction in clinical practice. Good documentation and understanding the extracted timestamps will prevent temporal leaks. Thorough verification of the extraction and preparation processes, data exploration and reproducible data cleaning can safeguard against data quality issues. Clinical assessment of the relevance and the sequence of extracted features and outcome within a patient admission by manual verification can further help avoiding problems (Tomašev et al., 2021). The solutions to specific issues can be implemented at either the extraction or preparation stage. Applicable to both data extraction and preparation are good understanding of the underlying data structure and current and historical healthcare processes, collaboration with relevant experts in conducting the work (Tomašev et al., 2021; Honeyford et al., 2022), and ensuring a qualitative process of extraction and preparation, supported by unit tests.
High-quality data extraction and preparation processes can support prediction models with clinical utility. The argument that EHR “data are not collected for research purposes” but for clinical use (Van Der Lei, 1991; Goldstein, 2020; Sauer et al., 2022; Maletzky et al., 2022; Thuraisingam et al., 2021; Arbet et al., 2021) has been made multiple times. Among various research goals, applied prediction models may be the most effective way to leverage electronic health record (EHR) data. Prediction models can exploit the comprehensive clinical information that is not always readily available to all healthcare professionals, such as nurses, doctors, hygienists, and other therapists, and have proven practical applicability (Lee et al., ).
@@
We hold the opinion that, in the context of prediction models, the extraction process should not attempt to correct errors residing in the EHR database in the attempt to align the data to the clinical reality and it should reflect the information from the EHR, presented in a simplified format. Extracting and preparing the training data in a different manner than for clinical implementation (e.g.: temporal leaks) poses the risk of potential overoptimistic evaluation, in the light of which the model’s performance when implemented in clinical practice will be deceiving. We acknowledge that there are divergent views on this topic and that in the context of inferential studies, when the analysis does not need to be reproduced on future data, corrections at extraction time might be preferred. At the same time, multi-purpose extractions (for both prediction and inferential studies) pose the challenge of solving the divergent view.
@@
Our work has several limitations. First, it is based on our experiences and a selective literature review. Insights from other research groups working with EHR data would likely highlight additional challenges and recommendations that we may have overlooked, offering a more exhaustive set of recommendations. Second, our focus was on single-site structured EHR data. We do not explicitly address combined data registries, such as national registries (integrating surveys, general practice, insurance data, and EHR extractions), multi-center data extractions, hospital data collected for clinical trials, or the extraction and processing of unstructured data (e.g., text notes or images). Extensions to multi-center datasets or unstructured data are possible. Third, we provide recommendations for problems that can impede the model implementation in practice, but we do not cover explicitly post-implementation challenges, although some of our recommendations can inform on monitoring checks that can be implemented, which remains a subject for medical device post-market surveillance. Fourth, while we emphasize the importance of high-quality data extraction and preparation as the foundation of reliable predictive models, we do not assess the impact of each challenge on the final prediction model. The impact will likely depend on the magnitude of problem, the subgroups in which it manifests, the prediction task or even the type of model (e.g., tree-based models are generally resilient to outliers). While the impact of dataset size, missing data and outcome definition on the model performance have been studied (van Os et al., 2022), the impact of other steps in the data preparation procedure remains unknown. Researchers identify and resolve problems during data preparation without assessing their influence on model performance, which we acknowledge having done the same. Finally, although we focused on models with clinical applicability, by recommending an implementation-ready process to avoid discrepancies between the training data and future data, we did not specifically address model implementation challenges and model generalization to new hospital settings. The requirements for reproducing the data extraction and preparation can vary significantly based on the EHR software in use (whether from the same or different vendor), the data extraction platform, and the specific hospital workflow. Sendak et al (Sendak et al., 2017) estimate “approximately 75% of the effort invested in the initial data preparation for developing prediction models must be reinvested for each hospital”. We believe that the estimation would vary widely, with standardized extractions from the same EHR software potentially requiring less time. Further research could focus on extending the list of challenges and recommendations based on the experience with EHR data of other research groups, extending the scope to larger extraction contexts or assessing the impact of erroneous or suboptimal extraction and preparation on the final model.
@@
The extensive list of challenges and practical recommendations for EHR data extraction and preparation presented here is intended for improving the quality of research and the practical applicability of clinical predictive models. Recognizing that there is no “one-size-fits-all” solution, implementing the strategies applicable to each specific prediction project can ultimately support robust, reproducible, and impactful healthcare predictions.

### v2

@@ 5 Discussion and Conclusion
In this work, we highlighted challenges and recommendations for the extraction and preparation of EHR data for predictive modeling. Adhering to the “garbage in, garbage out” principle, prediction models rely heavily on the quality and relevance of the input data to generate meaningful and reliable predictions. High-quality data extraction and preparation processes can support prediction models with clinical utility. While it is often argued that EHR “data are not collected for research purposes” (Van Der Lei, 1991; Goldstein, 2020; Sauer et al., 2022; Maletzky et al., 2022; Thuraisingam et al., 2021; Arbet et al., 2021), the fact that these data are sufficient for supporting the patient care suggests they could be adequate for developing prediction models. Even though the list of challenges may seem extensive, it should not discourage researchers new to working with EHR data. Not all challenges will apply to every project, and as EHR systems continue to evolve, many issues might affect only historical data and not current data at implementation time. Successfully leveraging EHR data requires both an understanding of its limitations and an appreciation of its potential. Despite its complexities, EHR data offer a rich, comprehensive source of real-world clinical information that can drive impactful research and improve patient outcomes. Applied prediction models can exploit the comprehensive clinical information that is not always readily available to all healthcare professionals, such as nurses, doctors, hygienists, and other therapists, and have proven practical applicability (Lee et al., 2020).
@@
Data extraction and preparation for predictive modelling using EHR data are resource-intensive processes, with time and cost varying depending on the maturity of the extraction framework, the prediction task and the team’s experience in working with EHR data. It is estimated that this phase (which generally includes collaboration with data extraction engineers and clinical experts), takes up to three months for an entire team (Tomašev et al., 2021), but such estimations will be highly dependent on the maturity of the extraction process and of the team’s experience with EHR data. Data quality often varies significantly across EHR systems and extraction processes, as noted by Weiskopf et al. (Weiskopf and Weng, 2013). Issues such as data gaps, temporal leaks, incorrect linking, inconsistencies in clinical concept and terminology mapping can affect the quality of extracted datasets and compromise the model’s performance and applicability. To address such challenges, we proposed a list of practical recommendations informed by our experiences with EHR data and insights from published studies. We organized the challenges and recommendations into: cohort definition, outcome definition, feature engineering and data cleaning; the first three categories can also be consulted when planning a project. A clear definition of the prediction task or research question, of the intended use and the intended users of a prediction model are a first critical step for defining the outcome, the cohort and the features of interest (Arbet et al., 2021; Honeyford et al., 2022). It is not uncommon to deem the EHR data inadequate for the prediction task, before proceeding with the model building phase (Thuraisingam et al., 2021).
@@
For cohort definition, we recommend extracting a broader patient context beyond the immediate focus, assessing the completeness of data used for inclusion/exclusion criteria and its availability at prediction time, preventing omissions or duplications, and careful definition of episodes of interest for prediction. Outcomes can be derived in different ways, each with advantages and shortcomings. We generally advise against the use of ICD codes (as these can be up- or undercoded and are generally not timestamped), unless carefully assessed as appropriate for the prediction task. Good understanding of hospital’s processes, thorough verification of outcomes derived in code, manual inspection of labels and agreement between data sources can also prevent incorrect outcome definition. Mapping terminology to coding systems (e.g., ICD, CPT, LOINC, SNOMED-CT or others) can facilitate feature engineering. However, not all items in the EHR system are aligned with standardized terminologies (e.g., LOINC codes for lab results are often not used). Implementing terminology mapping during the data extraction phase (e.g., through OMOP CDM) can significantly reduce the effort required during data preparation. High-quality data extraction documentation and good understanding of the underlying clinical concepts and healthcare processes are essential to support feature engineering. Data exploration can reveal unaddressed problems and further inform the construction of meaningful features. For feature engineering in prediction settings, the timestamp when a clinical item is available in the system is of greatest interest, as this is the time at which predictor values become available for prediction in clinical practice. Good documentation and correct interpretation the extracted timestamps will prevent temporal leaks. Thorough verification of the extraction and preparation processes (using manual and/or automated tests), data exploration and reproducible data cleaning can safeguard against data quality issues.
Clinical assessment of the relevance and the sequence of extracted features and outcome within a patient admission by manual verification of a random sample of admissions can further help detecting problems (Tomašev et al., 2021). The solutions to specific issues can be implemented at either the extraction or preparation stage. Applicable to both data extraction and preparation are good understanding of the underlying data structure and current and historical healthcare processes, collaboration with relevant experts in conducting the work (Tomašev et al., 2021; Honeyford et al., 2022), and ensuring a qualitative process of extraction and preparation, supported by unit tests. We recommend maintaining consistency in data extraction and preparation between model training and clinical implementation. Exceptions might though exist for correcting historical data problems that are not expected to reoccur in future data.
@@
We hold the opinion that, in the context of prediction models, the extraction process should not attempt to correct errors residing in the EHR database in the attempt to align the data to the clinical reality, but it should reflect the information from the EHR, presented in a simplified format. We recommend that data cleaning steps are performed during data preparation, in a pragmatic and programmatic manner that can be reproduced at implementation time. Extracting and preparing the training data in a different manner than for clinical implementation (e.g., temporal leaks) poses the risk of potential overoptimistic evaluation, in the light of which the model’s performance when implemented in clinical practice will be lower than expected. We acknowledge that there are divergent views on this topic and that in the context of inferential studies, when the analysis does not need to be reproduced on future data, corrections during data extraction might be preferred. At the same time, multi-purpose extractions (for both prediction and inferential studies) pose the challenge of solving this divergent view.
While the level of detail for reporting data extraction and preparation in published prediction studies varies, we do not provide specific recommendations on this aspect. Some researchers advocate for comprehensive documentation of these processes (de Kok et al., 2024; Chin et al., 2023; Hernandez-Boussard et al., 2020), others emphasize the importance of sharing the data preparation code to ensure transparency (Goldstein, 2024; Johnson et al., 2017). Data extraction for public datasets is generally documented in a separate publication. This can be complemented by sharing the data preparation code and describing in the main article the key differences between the raw extracted data and the final prepared dataset (de Kok et al., 2024). Each strategy has its advantages, depending on the audience, journal requirements, and the desired tradeoff between transparency and conciseness. While we advocate for transparent reporting, this article does not specifically address reporting guidelines for data extraction and preparation.
@@
Our work has several limitations. First, it is based on our experiences and a selective literature review that cannot be exhaustive. We acknowledge that every project will face use case specific challenges. Insights from other research groups working with EHR data would likely highlight additional challenges and recommendations that we may have overlooked, offering a more complete set of recommendations. Second, our focus was on single-site structured EHR data. Extensions to multi-center datasets or unstructured data are possible. Multi-center datasets pose additional challenges with regards to aligning clinical concepts and following the same patient in different hospitals and/or general practitioner systems. Patient care happens across multiple systems, and single-site extractions provide only a fragmented view of the entire patient care. Third, we provide recommendations for problems that can impede the model implementation in practice, but we do not cover explicitly post-implementation challenges, although some of our recommendations can inform on monitoring checks that can be implemented, which remains a subject for medical device post-market surveillance. Fourth, while we emphasize the importance of high-quality data extraction and preparation as the foundation of reliable predictive models, we do not assess the impact of each challenge on the final prediction model. The impact will likely depend on the magnitude of the problem, the subgroups in which it manifests, the prediction task or even the type of model (e.g., tree-based models are generally resilient to outliers). While the impact of dataset size, missing data and outcome definition on the model performance have been studied (van Os et al., 2022), the impact of other steps in the data preparation procedure remains unknown. Researchers identify and resolve problems during data preparation without assessing their impact on model performance, which we acknowledge having done the same. Research on measurement error (Luijken et al., 2020) demonstrates that inconsistencies in predictor definitions between training and test data can affect model calibration. A similar impact may occur if correction strategies differ between model training and clinical implementation. Finally, although we focused on models with clinical applicability, by recommending an implementation-ready process to avoid discrepancies between the training data and future data due to data extraction and preparation, we did not specifically address model generalization to new hospital settings. The requirements for reproducing the data extraction and preparation can vary significantly based on the EHR software in use (whether from the same or different vendor), the data extraction platform, and the hospital workflow and data registration procedures. Sendak et al (Sendak et al., 2017) estimate “approximately 75% of the effort invested in the initial data preparation for developing prediction models must be reinvested for each hospital”. We argue that the estimation would vary widely, with standardized extractions from the same EHR software potentially requiring less time. Further research could focus on extending the list of challenges and recommendations based on the experience with EHR data of other research groups, extending the scope to larger extraction contexts or assessing the impact of erroneous or suboptimal extraction and preparation on the final model.
@@
The extensive list of challenges and practical recommendations for EHR data extraction and preparation presented here is intended for improving the quality of research and the practical applicability of clinical predictive models. Since all modeling efforts begin with the underlying data, failing to address data quality issues risks producing unreliable and non-generalizable models. Our focus extends beyond the initial data curation stage in the AI life cycle (Ng et al., 2022); we also address early-stage issues that can ultimately negatively impact model deployment. Recognizing that there is no “one-size-fits-all” solution, our list of challenges and recommendations, though not exhaustive, is comprehensive enough to support many EHR-based prediction projects. Implementing the strategies applicable to each project can ultimately enhance robustness, reproducibility, and real-world impact of EHR-based prediction models.
6 Abbreviations
AKI: Acute Kidney Injury
API: Application Programming Interface
AVPU: Alert, Voice, Pain, Unresponsive
CAM: Confusion Assessment Method
CCS: Clinical Classifications Software
CCSR: Clinical Classifications Software Refined
CDC: Centers for Disease Control and Prevention
CLABSI: Central Line-Associated Bloodstream Infection
COVID-19: Coronavirus Disease 2019
CPT: Current Procedural Terminology
DRS: Delirium Rating Scale
ETL: Extract, Transform, Load
EHR: Electronic Health Record
FiO2: Fraction of Inspired Oxygen
HIS: Hospital Information System
ICD: International Classification of Diseases
ICU: Intensive Care Unit
NA: Not Applicable
OMOP CDM: Observational Medical Outcomes Partnership Common Data Model
PEEP: Positive End-Expiratory Pressure
VAE: Ventilator-Associated Event